Herman’s Hot Chocolate stand sold 12 cups of hot chocolate and made a profit of $16. The next day they sold 10 cups of hot chocolate and made a profit of $12.50. How much did they charge for each cup of hot chocolate?

a)$1.25 b) $1.33 c) $1.50 d) $1.75 e) $1.85

Answer by richwmiller(17219)

12*x=16
x=16/12=4/3=1.25 profit per cup
10*x=12.50
x=1.25 profit per cup
Not enough info to know how much was charged .
We are only told the profit and not the costs nor revenue.
Profit is not the same as revenue (charged).